Given P(E) = 0.25, P(F) = 0.6, and P(E ∪ F) = 0.7.
Find:
1
Expert's answer
2020-08-03T18:12:53-0400
"P(E\cup F)=P(E)+P(F)-P(E\cap F)"

"P(E\cap F)=P(E)+P(F)-P(E\cup F)="

"=0.25+0.6-0.7=0.15"

Since "P(E\cap F)\not =0," then event "E" and event "F" are not mutually exclusive.



"P(E)\cdot P(F)=0.25\cdot0.6=0.15=P(E\cap F)"

Event "E" and event "F" are independent.